OpenStudy (babynini):

Simplifying: tan(theta) + cos (-theta) + tan(-theta) how do I get rid of negative thetas?

OpenStudy (babynini):

\[\tan (\Theta)+\cos(-\Theta)+\tan(-\Theta)\] \[(\sin \Theta/\cos \Theta) + \cos(-\Theta) + (\sin/\cos)(-\Theta) \]

OpenStudy (babynini):

@optiquest :)

OpenStudy (anonymous):

repost in math

OpenStudy (babynini):

ah! sorry! thanks haha